OpenStudy (anonymous):

could some one please explain how to solve this problem graphically? 2x − y = 4, 5x − y = 13 i know i need to put it into y=mx+b form, but I am not sure if i did ti correct because i am not used to having a negative y. I got y=-2x-4 and y=-5x-13 is this correct and if so how would I graph it? please help graphing is such a weak point for me

OpenStudy (anonymous):

Start by mulitplying both sides by -1 Then you don't have a -y

OpenStudy (anonymous):

How is the 5x positive? and the 2x and 4 positive? when I subtract them they become negative i thought

OpenStudy (anonymous):

I am completely confused now

OpenStudy (anonymous):

If you multiply both side of the equation by -1 then you just change all the signs before you start to move things around: -2x + y = -4

OpenStudy (anonymous):

oh ok I see that is how it becomes y=2x-4 and y=5x-13 is this correct?

OpenStudy (anonymous):

But he's trying to put into this form: y = mx + b So it won't help to move the y

OpenStudy (anonymous):

im a she lol

OpenStudy (anonymous):

please please help me get this. i am in the last of the last of this class my final is thursday I gotta get it

OpenStudy (anonymous):

k ty

OpenStudy (anonymous):

Omg I get it thank you so much